Find the tension in cables B and C if w1 = 73.0 N and w2 = 170 N.

60°у 60°
Answer #1

60° 60° TB Tc TA

Applying Newton’s 2nd law horizontally at the knot,

Fnet = ma

TCx -TBx = 0

TCx =TBx

TCcos60=TBcos60

TC =TB--------------(1)

Applying Newton’s 2nd law vertically at the knot,

Fnet = ma

TCy +TBy –TAy = 0

TCsin60+TBsin60 = TAy ……………(2)

Plug (1) in (2)

TBsin60+TBsin60 = TAy

2TBsin60 = TAy……………(3)

Applying Newton’s 2nd law vertically at the knot,

Fnet = ma

TAy-W1-W2 = 0

TAy=W1+W2---------------(4)

Plug (4) in (3),

2TBsin60 = W1+W2

TB = (W1+W2)/(2sin60)

Plug values,

TB =(73.0+170)/(2sin60)

TB = 140.3 N

From (1),

TC =TB = 140.3 N